# QWHA: Quantization-Aware Walsh-Hadamard Adaptation for Parameter-Efficient Fine-Tuning on Large Language Models

### v1

@@ 3.2 Quantization-Aware Adapter Initialization
where $\lVert\cdot\rVert_{F}$ denotes Frobenius norm. Following the reduction procedure used in frantar2023gptq and deng2025cloq, this reduces to:
@@
Here, ${\bm{R}}={\bm{U}}\mathbf{\Sigma}^{1/2}$ is the invertible square root of the Hessian matrix attained by SVD as ${\bm{X}}{\bm{X}}^{\top}={\bm{U}}\mathbf{\Sigma}{\bm{U}}^{\top}$. A detailed derivation on this reduction is provided in Appendix C.1. As we aim to find a sparse ${\bm{F}}({\bm{c}},{\bm{E}})$ that minimizes Equation 6, it constitutes an NP-hard sparse approximation problem (SAP) (natarajan1995). To make this problem more tractable, we decompose it into two subproblems: first, parameter selection to determine the locations of the nonzero elements to fine-tune ($\bm{E}$}); and second, value refinement to optimize the values of the selected positions ($\bm{c}$}).
@@
Figure 4: Rank of adapter weights.
@@
Figure 5: Layer output error.
This allows the selected basis vectors to account for the impact of unselected vectors, yielding a more accurate approximation. Without this step, interactions among basis vectors are ignored, leading to suboptimal error reduction. Note that the refinement is applicable regardless of the parameter selection strategy. Figure 5 shows that refinement is crucial for reducing layer output error, presenting the averaged layer output error of a model with parameters selected by AdaAlloc. Further details on the error analysis are provided in Appendix C.3. Finally, $\bm{E}$} and $\bm{c}$} for channel $i$ are initialized to the selected indices and their refined values $\bm{x}$$^{*}$. Algorithm 1 summarizes the initialization process, with details provided in Appendix C.4.
@@ 4 Experiments
We evaluate QWHA on the Mistral-7B-v0.3 (mistralai2024mistral7bv03) and LLaMA (grattafiori2024llama) model families, including LLaMA-3.1-8B and LLaMA-3.2-3B. For instruction fine-tuning, we use the Stanford-Alpaca dataset (alpaca) [^1] https://huggingface.co/datasets/yahma/alpaca-cleaned with 52k samples. We evaluate on zero-shot commonsense question answering (CSQA)(eval-harness), covering seven multiple-choice benchmarks(clark2018think; clark2019boolq; zellers2019hellaswag; talmor2019commonsenseqa; bisk2020piqa; sakaguchi2021winogrande). For arithmetic reasoning, we adopt the GSM8k (gsm8k) dataset and evaluate zero-shot accuracy following gsm8k.
@@
We include full fine-tuned model (Fine-tuned) and quantized model, which use GPTQ (frantar2023gptq) with MagR (zhang2024magrweightmagnitudereduction) ($\mathrm{GPTQ_{MagR}}$) as baselines. We note that our method is also compatible with any other quantization schemes. We also include CLoQ, a recent QA-PEFT method that shares our goal of layer output error reduction during initialization for low-rank adapters. Other LoRA-based methods (kim2024ra; liao2024apiq) involving layer-wise calibration or layer-wise parameter allocation are orthogonal to our approach and can be integrated in future work. We evaluate sparse adapters, including SSH and LoCA (FT-based) and SHiRA (non FT-based). We note that LoCA further fine-tunes the randomly selected parameter indices via reparameterization with a cost of additional training overhead. We also build advanced hybrid baselines that integrate transforms or parameter selection strategies from prior works into our schemes by applying DCA and DHA with our AdaAlloc, or applying various parameter selection strategies to our WHA.
@@
Following prior work, adapters are applied to linear layers with a parameter budget of $P(r=64)$, and quantization is performed with a group size of 64. Note that we apply a constant scaling factor $\alpha$ to all adapters, while the equations in the preceding sections omitted it by $\alpha=1$ for simplicity. We set the AdaAlloc temperature to $t=1$, which suffices to meet the full-rank condition. We use WikiText-2 (wiki2016merity) as a calibration dataset for adapter initialization, following deng2025cloq, to ensure generality. All experiments are conducted on NVIDIA A100 80GB GPUs, and training hyperparameters are provided in Appendix D.1.
@@ 4.1 Fine-tuned Model Accuracy
We further examine the effectiveness of WHA and AdaAlloc, with QWHA consistently outperforming both low-rank adapters and advanced variants of sparse adapters. Figure LABEL:fig:rank_ablation for 4-bit quantized LLaMA-3.2-3B shows that increasing the number of parameters in CLoQ cannot close the accuracy gap with QWHA, as QWHA with $P(r>32)$ already surpasses CLoQ’s maximum achievable score. This highlights the advantage of WHA, which provides superior representational capacity than low-rank adapters. Table LABEL:table:eval_detail further demonstrates that WHA and AdaAlloc achieve the best results in each respective category of adapter type and parameter selection method. We note that LoCA’s post-hoc location selection undermines the effectiveness of quantization-aware initialization based on the initially chosen parameters, unlike in PEFT. Additional ablations on quantization group size are provided in Appendix D.3.
@@ C.1 Objective Function
We reduce the original optimization objective in Equation 5 to the form in Equation 6, following the approach of frantar2023gptq; deng2025cloq. Using the notations from Section 3.2, the reduction proceeds as follows:
@@
where ${\bm{R}}={\bm{U}}\mathbf{\Sigma}^{1/2}\in\mathbb{R}^{d_{\text{in}}\times d_{\text{in}}}$ is an invertible square root of the Hessian Gram matrix ${\bm{X}}{\bm{X}}^{\top}$, such that ${\bm{X}}{\bm{X}}^{\top}={\bm{U}}\mathbf{\Sigma}{\bm{U}}^{\top}$. Following deng2025cloq, we add a small regularization term $\lambda=0.0001\cdot\operatorname{\mathrm{tr}}({\bm{X}}{\bm{X}}^{\top})/d_{\text{in}}$ to the diagonal if $\bm{R}$} is not originally invertible. This reduction allows us to replace $\bm{X}$} with $\bm{R}$}, enabling efficient and effective calibration using multiple input data points. Rather than solving the optimization problem separately for each sample $\bm{X}$}, we can accumulate the contribution of activations via $\bm{R}$} and solve a single reduced problem.

### v3

@@ 3.2 Quantization-Aware Adapter Initialization
where $\lVert\cdot\rVert_{F}$ denotes Frobenius norm. Following the reduction procedure used in Frantar et al. (2023) and Deng et al. (2025), this reduces to:
@@
Here, ${\bm{R}}={\bm{U}}\mathbf{\Sigma}^{1/2}$ is the invertible square root of the Hessian matrix attained by SVD as ${\bm{X}}{\bm{X}}^{\top}={\bm{U}}\mathbf{\Sigma}{\bm{U}}^{\top}$. A detailed derivation on this reduction is provided in Appendix C.1. As we aim to find a sparse ${\bm{F}}({\bm{c}},{\bm{E}})$ that minimizes Equation 6, it constitutes an NP-hard sparse approximation problem (SAP) (Natarajan, 1995). To make this problem more tractable, we decompose it into two subproblems: first, parameter selection to determine the locations of the nonzero elements to fine-tune ($\bm{E}$}); and second, value refinement to optimize the values of the selected positions ($\bm{c}$}).
@@
Figure 4: Rank of adapter weights for each parameter selection methods.
@@
Figure 5: Effect of refinement on average layer output error.
This allows the selected basis vectors to account for the impact of unselected vectors, yielding a more accurate approximation. Without this step, interactions among basis vectors are ignored, leading to suboptimal error reduction. Note that the refinement is applicable regardless of the parameter selection strategy. Figure 5 shows that refinement is crucial for reducing layer output error, presenting the layer output error after initialization with parameters selected by AdaAlloc. Further details on the error analysis are provided in Appendix C.3. Finally, $\bm{E}$} and $\bm{c}$} for channel $i$ are initialized to the selected indices and their refined values $\bm{x}$$^{*}$. Algorithm 1 summarizes the initialization process, with details provided in Appendix C.4.
@@ 4 Experiments
We evaluate QWHA on the Mistral-7B-v0.3 (Mistral AI, 2024) and LLaMA (Grattafiori et al., 2024) model families, including LLaMA-3.1-8B and LLaMA-3.2-3B. For instruction fine-tuning, we use the Stanford-Alpaca dataset (Taori et al., 2023) [^1] https://huggingface.co/datasets/yahma/alpaca-cleaned with 52k samples. We evaluate on zero-shot commonsense question answering (CSQA)(Gao et al., 2024a), covering seven multiple-choice benchmarks(Clark et al., 2018; 2019; Zellers et al., 2019; Talmor et al., 2019; Bisk et al., 2020; Sakaguchi et al., 2021). For arithmetic reasoning, we adopt the GSM8k (Cobbe et al., 2021) dataset and evaluate zero-shot accuracy following Cobbe et al. (2021).
@@
We include full fine-tuned model (Fine-tuned) and quantized model, which use GPTQ (Frantar et al., 2023) with MagR (Zhang et al., 2024) ($\mathrm{GPTQ_{MagR}}$) as baselines. We note that our method is also compatible with any other quantization schemes. We also include CLoQ, a recent QA-PEFT method that shares our goal of layer output error reduction during initialization for low-rank adapters. Other LoRA-based methods (Kim et al., 2024a; Liao et al., 2024) involving layer-wise calibration or layer-wise parameter allocation are orthogonal to our approach and can be integrated in future work. We evaluate sparse adapters, including SSH and LoCA (FT-based) and SHiRA (non FT-based). We note that LoCA further fine-tunes the randomly selected parameter indices via reparameterization with a cost of additional training overhead. We also build advanced hybrid baselines that integrate transforms or parameter selection strategies from prior works into our schemes by applying DCA and DHA with our AdaAlloc, or applying various parameter selection strategies to our WHA.
@@
Following prior work, adapters are applied to linear layers with a parameter budget of $P(r=64)$, and quantization is performed with a group size of 64. Note that we apply a constant scaling factor $\alpha$ to all adapters, while the equations in the preceding sections omitted it by $\alpha=1$ for simplicity. We set the AdaAlloc temperature to $t=1$, which suffices to meet the full-rank condition. We use WikiText-2 (Merity et al., 2016) as a calibration dataset for adapter initialization, following Deng et al. (2025), to ensure generality. All experiments are conducted on NVIDIA A100 80GB GPUs, and training hyperparameters are provided in Appendix D.1.
@@ 4.1 Fine-tuned Model Accuracy
We further examine the effectiveness of WHA and AdaAlloc, with QWHA consistently outperforming both low-rank adapters and advanced variants of sparse adapters. Figure 6 for 4-bit quantized LLaMA-3.2-3B shows that increasing the number of parameters in CLoQ cannot close the accuracy gap with QWHA, as QWHA with $P(r>32)$ already surpasses CLoQ’s maximum achievable score. This highlights the advantage of WHA, which provides superior representational capacity than low-rank adapters. Table 4.1 further demonstrates that WHA and AdaAlloc achieve the best results in each respective category of adapter type and parameter selection method. We note that LoCA’s post-hoc location selection undermines the effectiveness of quantization-aware initialization based on the initially chosen parameters, unlike in PEFT. Additional ablations on quantization group size are provided in Appendix D.3.
Table 4: Accuracy (%) evaluation results on CSQA and GSM8k benchmarks with variants of adapter types and parameter selection strategies in LLaMA-3.2-3B. ‘QA Init.’ denotes the existence of quantization-aware initialization, and ‘Refine.’ denotes the value refinement during initialization.
Adapter QA Coefficient Refine. 4-bit 3-bit 2-bit
Type Init. Selection CSQA GSM8k CSQA GSM8k CSQA GSM8k
WHA Random 66.00 40.94 63.53 37.60 54.03 24.41
WHA Random 65.91 40.71 63.91 37.30 54.48 24.48
WHA Magnitude 66.07 41.01 64.52 36.69 56.49 28.12
WHA LoCA 65.75 40.94 63.73 36.92 53.93 21.15
WHA SSH 65.96 40.78 62.92 36.92 54.20 27.14
WHA AdaAlloc 66.11 41.47 64.80 39.58 57.03 29.11
gray [1pt/1pt] DCA AdaAlloc 65.54 39.72 64.77 37.30 55.95 27.29
DHA AdaAlloc 65.92 40.84 64.35 38.89 56.05 27.52
Sparse AdaAlloc 65.60 40.94 63.43 37.53 55.97 26.54
black
Figure 6: Accuracy of CLoQ and QWHA.
Table 5: Training time (hours) on Alpaca dataset.
Batch CLoQ SHiRA QWHA SSH LoCA
Size
1 12.5 15.5 18.2 63.3 92.3
2 7.1 8.2 9.7 45.8 53.4
4 5.0 5.5 6.0 26.1 30.1
8 4.1 4.3 4.6 13.3 16.5
16 3.6 3.7 3.9 8.3 9.8
4.2 Computational Efficiency
Training time for QWHA on the Alpaca dataset with LLaMA-3.1-8B is reported in Table 6, where QWHA achieves a substantial speedup over previous FT-based adapters by leveraging WHA. WHA employs a single transform instead of the double transform used in conventional FT-based adapters, while achieving higher accuracy. Moreover, the fast recursive WHT kernel replaces matrix multiplications with a smaller number of additions and subtractions. In contrast, the recursive kernels of DCT and DHT, which require the DFT, are slower than direct matrix multiplication due to duplicated computations for imaginary parts. As a result, WHT achieves a similar training time to low-rank adapters or to a simple sparse adapter. In contrast, LoCA incurs additional latency even compared to SSH, due to the training of location parameters. Memory usage remains almost identical across all adapters with the same number of parameters. Detailed results on the training time of each transform kernel and the memory usage of each adapter are provided in Appendix D.4 and Appendix D.5.
5 Conclusion
In this work, we introduce QWHA, a novel QA-PEFT framework featuring a Walsh-Hadamard transform-based adapter and its quantization-aware parameter initialization scheme. WHA offers strong fine-tuning capability and excels in quantization error reduction. The proposed AdaAlloc scheme facilitates both fine-tuning and quantization error reduction during parameter selection, while parameter refinement enables substantial quantization error reduction. We validate QWHA across diverse models and datasets, where it consistently outperforms existing baselines in accuracy and demonstrates its effectiveness. We also show that using WHA with a single transform provides computational benefits, enabling more efficient training than conventional FT-based adapters.
Appendix A Quantization Error Distribution
We present the distribution of quantization errors and their relationship to outliers in the pre-trained weights, as discussed in Section 2.1, in Figure 7. Figure 7(a) shows the overall error distribution, while Figure 7(b) highlights the channel-wise similarity between quantization errors and pre-trained weights in the 14th layer of LLaMA-3.2-3B. During quantization, values are divided by the quantization scale, typically defined per group within each output channel, and then rounded to an integer and clamped within a range determined by the bit-width. Most quantization errors remain within this rounding range, but large-magnitude outliers are often clamped, leading to large errors. Because model accuracy is highly sensitive to outlier weights, their quantization errors can significantly degrade performance. In QA-PEFT, it is therefore crucial to mitigate such outlier-induced errors during initialization by adapting the weights, particularly for large-magnitude values originating from salient outliers.
Figure 7: (a) Weight quantization error distribution and (b) its channel-wise similarity to the pre-trained weights in 14th layer Key projection of 4-bit quantized LLaMA-3.2-3B. In Figure (b), each pixel represents the $\ell_{2}$ norm of weight quantization errors (left) and that of pre-trained weights (right) for each output channel ordered by channel index from top-left to bottom-right.
Appendix B WHT-based Adapter (WHA)
B.1 FT-based Adapter Kernels
We describe a class of Fourier-related transform (FT) kernels employed in our adapters and prior studies in this section (Gao et al., 2024b; Du et al., 2025; Shen et al., 2025).
Walsh-Hadamard Transform (WHT).
The Walsh-Hadamard Transform (WHT) matrix $\bm{H}$} introduced in Equation 4 is constructed following conventions in prior works (Tseng et al., 2024; Ashkboos et al., 2024). For a dimension $N=2^{n}$, a WHT matrix ${\bm{H}}\in\mathbb{R}^{N\times N}$ is defined recursively as the Kronecker product of smaller matrices:
${\bm{H}}_{2}=\frac{1}{\sqrt{2}}\begin{bmatrix}1&1\\ 1&-1\end{bmatrix},\quad{\bm{H}}_{N}={\bm{H}}_{2}\otimes{\bm{H}}_{2^{n-1}},$ (10)
where $\otimes$ denotes the Kronecker product. For non-power-of-two dimensions, Hadamard matrices exist for certain values (Seberry & Yamada, 1992; Hedayat et al., 1999; Gerakoulis & Ghassemzadeh, 2004), which can be retrieved from Sloane (2004). More generally, for $N=2^{n}\cdot m$, where ${\bm{H}}_{m}$ is a known Hadamard matrix, the transform is defined as:
${\bm{H}}_{N}={\bm{H}}_{2^{n}}\otimes{\bm{H}}_{m}.$ (11)
The rows of ${\bm{H}}_{N}$ form an orthogonal basis, known as Walsh-Hadamard bases, satisfying:
${\bm{H}}_{N}^{\top}{\bm{H}}_{N}={\bm{H}}_{N}{\bm{H}}_{N}^{\top}={\bm{I}}_{N}.$ (12)
The matrix ${\bm{H}}_{2^{n}}$ can be computed in $O(n\log n)$ time (Kunz, 1979). In practice, ${\bm{H}}_{N}$ can be precomputed once and cached for reuse across layers of the same size, incurring negligible cost in both computation and memory. To further accelerate computation, we employ the Fast Hadamard multiplication kernel from Dao-AILab (2024), which avoids explicit matrix construction by using a fused kernel of only additions and subtractions.
Discrete Fourier Transform (DFT).
FourierFT (Gao et al., 2024b) was the first study of FT-based adapters and used the discrete Fourier transform (DFT). The transform kernel ${\bm{H}}\in\mathbb{C}^{N\times N}$ is defined as:
$H_{jk}=\frac{1}{\sqrt{N}}e^{-i\frac{2\pi jk}{N}}=\frac{1}{\sqrt{N}}\left\{\cos\left(\frac{2\pi jk}{N}\right)-i\sin\left(\frac{2\pi jk}{N}\right)\right\},\quad 0\leq j,k<N.$ (13)
Although effective, later works adopted real-valued FT variants to avoid the complex-domain nature of the DFT, since deep learning frameworks typically discard the imaginary components and compute only with the real values.
Discrete Hartley Transform (DHT).
SSH (Shen et al., 2025) employs the discrete Hartley transform (DHT), a real-valued variant of the FT with kernel:
$H_{jk}=\Re\big(\frac{1}{\sqrt{N}}e^{-i\frac{2\pi jk}{N}}\big)-\Im\big(\frac{1}{\sqrt{N}}e^{-i\frac{2\pi jk}{N}}\big)=\frac{1}{\sqrt{N}}\text{cas}\left(\frac{2\pi jk}{N}\right),\quad 0\leq j,k<N,$ (14)
where $\text{cas}(x)=\cos x+\sin x$.
Discrete Cosine Transform (DCT).
LoCA (Du et al., 2025) employs another real-valued FT, the discrete cosine transform (DCT), whose kernel is:
$H_{jk}=\begin{cases}\frac{1}{\sqrt{N}}&j=0\\ \sqrt{\frac{2}{N}}\cos\left(\frac{\pi(2k+1)j}{2N}\right)&0<j<N\end{cases}\quad,\;0\leq k<N.$ (15)
B.2 Rank of WHA
This section provides a detailed explanation of the full-rank property of WHA and its conditions, as discussed in Section 3.1 and illustrated in Figure 2(a). To preserve the expressiveness of a fine-tuned model under a limited parameter budget, it is critical to ensure high rank capacity in the weight update. Unlike low-rank adapters, which inherently restrict the parameter subspace, WHA is sparsely structured yet can retain high representational capacity by maintaining full rank. This also holds in typical sparse adapters, including FT-based adapters.
We build on theoretical insights from prior work on sparse random matrices Coja-Oghlan et al. (2020), which provides conditions under which such matrices are full rank. Specifically, consider a random sparse matrix ${\bm{F}}\in\mathbb{R}^{d_{\text{out}}\times d_{\text{in}}}$, where each input and output channel has $k$ and $l$ non-zero entries on average. Then, $\bm{F}$} is full rank when $k,l\geq 2$ as $d_{\text{in}},d_{\text{out}}\to\infty$, and thus full rank with high probability. Following the notations in Coja-Oghlan et al. (2020), we derive the corresponding condition for our setting to guarantee full-rank behavior in WHA.
Condition Function.
We define the probability generating functions for the distributions of random non-zero entries per column and per channel. Given that these distributions are degenerate, the generating functions and their derivatives are:
$D(z)=z^{k},\quad D^{\prime}(z)=kz^{k-1},\quad D^{\prime}(1)=k,$ (16)
$K(z)=z^{l},\quad K^{\prime}(z)=lz^{l-1},\quad K^{\prime}(1)=l,$ (17)
Then, the condition function $\Phi(z)$ that determines the full rank condition is given by:
$\Phi(z)=D\left(1-\frac{K^{\prime}(z)}{l}\right)-\frac{k}{l}\left[1-K(z)-(1-z)K^{\prime}(z)\right].$ (18)
To ensure the full rank of the matrix $A$, the inequality must hold as:
$\Phi(z)<\Phi(0),\quad\forall\,0<z\leq 1,$ (19)
Substituting the explicit forms for $D(z),K(z),D^{\prime}(z),K^{\prime}(z)$ into Equation 18 yields the right hand side as:
$\Phi(z)=(1-z^{l-1})^{k}-\frac{k}{l}+kz^{l-1}-\frac{k(l-1)}{l}z^{l}.$ (20)
As $\Phi(z=0)=1-\frac{k}{l}$, the condition in Equation 19 finally simplifies to:
$(1-z^{l-1})^{k}+kz^{l-1}-\frac{k(l-1)}{l}z^{l}-1<0,\quad 0<z\leq 1.$ (21)
Practical Considerations.
The inequality in Equation 21 shows that the condition generally holds for integers $k,l\geq 2$. For the total number of parameters $p=r(d_{\text{in}}+d_{\text{out}})$ with $r\geq 2$, we have $k=p/d_{\text{in}}>r$ and $l=p/d_{\text{out}}>r$ under random selection, thus satisfying the full-rank condition when $d_{\text{in}},d_{\text{out}}$ are sufficiently large. Importantly, since AdaAlloc selects indices independently within each channel, the parameters across input rows can be treated as uniformly random selections. Moreover, AdaAlloc’s per-channel allocation with remainder assignment and temperature control guarantees at least two elements in every channel (i.e., $l\geq 2$). Empirically, parameter budgets corresponding to $P(r\geq 4)$ ensure at least two elements per row (i.e., $k\geq 2$), even for linear layers with large output dimensions, which might otherwise receive few parameters per input row. Hence, the full-rank conditions hold, and the matrix $\bm{F}$} in QWHA is nearly full rank.
B.3 Energy Concentration of WHT
In this section, we quantify the energy concentration property of WHT discussed in Section 3.1, using Figure 2(b) and Figure 3(a).
Distribution of Singular Values and Coefficients.
Figure 8 presents the distributions of singular values from SVD and of transform coefficients sorted by their squared magnitudes. Here, the area under each plot is equal to $\lVert\Delta{\bm{W}}_{Q}\rVert_{F}^{2}$ (details in the following paragraph). The distributions follow a Pareto-like behavior, where sharpness can be quantified using the hill index $\eta$. The Pareto hill index is a value which implies the heaviness of a tail. In fact, this is the reciprocal of the Pareto tail index, defined as the mean of the log ratios of consecutive order statistics from the top-k largest magnitudes. A smaller hill index $\eta$ implies faster convergence of the cumulative distribution (Arnold, 1983). Hence, WHT exhibits the sharpest distribution, making it feasible to retain more information with fewer parameters $P(r)$ when implemented in a sparse adapter, as shown in Figure 2.
Figure 8: Singular value and coefficient magnitude (squared) distributions with the Pareto hill index $\eta$ in the 14th-layer Key projection of LLaMA-3.2-3B.
Energy of Singular Values and Coefficients.
Throughout this work, we use the term energy to denote the squared $\ell_{2}$ norm of the singular values in a decomposition or the spectral coefficients in a transform. We show that the total energy of both SVD and orthonormal transforms reduces to the Frobenius norm of the transformed matrix. (For example, in Figure 8, the area under each curve corresponds to $\lVert\Delta{\bm{W}}_{Q}\rVert_{F}^{2}$.)
Proposition 1.
Let ${\bm{M}}\in\mathbb{R}^{m\times n}$ be a matrix, and let its singular value decomposition (SVD) be ${\bm{M}}={\bm{U}}\mathbf{\Sigma}{\bm{V}}^{\top}$, where ${\bm{U}}\in\mathbb{R}^{m\times m}$ and ${\bm{V}}\in\mathbb{R}^{n\times n}$ are orthonormal matrices, and $\mathbf{\Sigma}\in\mathbb{R}^{m\times n}$ is a diagonal matrix with entries $\sigma_{i}$ on the diagonal for $i=1,\ldots,\min(m,n)$. Define ${\bm{F}}={\bm{M}}{\bm{H}}$ for an orthonormal matrix ${\bm{H}}\in\mathbb{R}^{n\times n}$. Then, we have:
$\lVert{\bm{M}}\rVert_{F}^{2}=\sum_{i=1}^{\min(m,n)}\sigma_{i}^{2}=\lVert{\bm{F}}\rVert_{F}^{2}.$
Proof.
@@ C.1 Objective Function
We reduce the original optimization objective in Equation 5 to the form in Equation 6, following the approach of Frantar et al. (2023); Deng et al. (2025). Using the notations from Section 3.2, the reduction proceeds as follows:
@@
where ${\bm{R}}={\bm{U}}\mathbf{\Sigma}^{1/2}\in\mathbb{R}^{d_{\text{in}}\times d_{\text{in}}}$ is an invertible square root of the Hessian Gram matrix ${\bm{X}}{\bm{X}}^{\top}$, such that ${\bm{X}}{\bm{X}}^{\top}={\bm{U}}\mathbf{\Sigma}{\bm{U}}^{\top}$. Following Deng et al. (2025), we add a small regularization term $\lambda=0.0001\cdot\operatorname{\mathrm{tr}}({\bm{X}}{\bm{X}}^{\top})/d_{\text{in}}$ to the diagonal if $\bm{R}$} is not originally invertible. This reduction allows us to replace $\bm{X}$} with $\bm{R}$}, enabling efficient and effective calibration using multiple input data points. Rather than solving the optimization problem separately for each sample $\bm{X}$}, we can accumulate the contribution of activations via $\bm{R}$} and solve a single reduced problem.